OpenStudy (anonymous):

The equation of the circle whose center is at (4,4) and whose radius is 5 is (x + 4)² + (y + 4)² = 5 (x + 4)² + (y + 4)² = 25 (x - 4)² + (y - 4)² = 25

OpenStudy (anonymous):

the equation for a circle is (x-a)^2 +(y-b)^2= r^2 look at the last one the radius is 5 so 5^2=25 so that checks look at a and b those are the centre points looks like (4,4) fits